inputenc.sty

Alan Jeffrey  Frank Mittelbach

v0.992 1999/09/17

1 Introduction

This package allows the user to specify an input encoding (for example, ASCII, ISO Latin-1 or Macintosh) by saying:

  \usepackage[encoding name]{inputenc}

The encoding can also be selected in the document with:

  \inputencoding{encoding name}

The only practical use of this command within a document is when using text from several documents to build up a composite work such as a volume of journal articles. Therefore this command will be used only in vertical mode.

The encodings provided by this package are:

• ascii ASCII encoding for the range 32–127.

• latin1 ISO Latin-1 encoding.

• latin2 ISO Latin-2 encoding.

• latin3 ISO Latin-3 encoding.

• latin5 ISO Latin-5 encoding.

• decmulti DEC Multinational Character Set encoding.

• cp850 IBM 850 code page.

• cp852 IBM 852 code page.

• cp437 IBM 437 code page.

• cp437de IBM 437 code page (German version).

• cp865 IBM 865 code page.

• applemac Macintosh encoding.

• next Next encoding.

• ansinew Windows 3.1 ANSI encoding, extension of Latin-1.

• cp1252 Synonym for ansinew.

• cp1250 Windows 1250 (central and eastern Europe) code page.

Each encoding has an associated .def file, for example latin1.def which defines the behaviour of each input character, using the commands:

     \DeclareInputText{slot}{text}
     \DeclareInputMath{slot}{math}

This defines the input character slot to be the text material or math material respectively. For example, latin1.def defines slots "D6 (Æ) and "B5 (μ) by saying:

    \DeclareInputText{214}{\AE}
    \DeclareInputMath{181}{\mu}

Note that the commands should be robust, and should not be dependent on the output encoding. The same slot should not have both a text and a math declaration for it. (This restriction may be removed in future releases of inputenc).

The .def file may also define commands using the declarations:

\providecommand or \ProvideTextCommandDefault. For example:
    \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}
    \DeclareInputText{188}{\textonequarter}

The use of the ‘provide’ forms here will ensure that a better definition will not be over-written; their use is recommended since, in general, the best definition depends on the fonts available.

See the documentation in fntguide.tex and ltoutenc.dtx for details of how to declare text commands.

2 Announcing the files

We announce the files:

 1 ⟨package⟩\NeedsTeXFormat{LaTeX2e}[1995/12/01]
 2 ⟨package⟩\ProvidesPackage{inputenc}
 3 ⟨ascii⟩ \ProvidesFile{ascii.def}
 4 ⟨latin1⟩ \ProvidesFile{latin1.def}
 5 ⟨latin2⟩ \ProvidesFile{latin2.def}
 6 ⟨latin3⟩ \ProvidesFile{latin3.def}
 7 ⟨latin5⟩ \ProvidesFile{latin5.def}
 8 ⟨decmulti⟩ \ProvidesFile{decmulti.def}
 9 ⟨cp850⟩ \ProvidesFile{cp850.def}
10 ⟨cp852⟩ \ProvidesFile{cp852.def}
11 ⟨cp437⟩ \ProvidesFile{cp437.def}
12 ⟨cp437de⟩ \ProvidesFile{cp437de.def}
13 ⟨cp865⟩ \ProvidesFile{cp865.def}
14 ⟨applemac⟩ \ProvidesFile{applemac.def}
15 ⟨next⟩ \ProvidesFile{next.def}
16 ⟨ansinew⟩ \ProvidesFile{ansinew.def}
17 ⟨cp1252⟩ \ProvidesFile{cp1252.def}
18 ⟨cp1250⟩ \ProvidesFile{cp1250.def}
19 [1999/09/17 v0.992 Input encoding file ]

Now we insert a \makeatletter at the beginning of all .def files.

20 ⟨−package⟩\makeatletter

3 The package

Before we start with the code, an important comment is in order: as you may or may not know, the tabbing environment changes the definition of the commands \', \`, and \=. Outside such an environment these commands produce the corresponding accents, inside they are used for special text positioning, and the accents can be accessed using \a', \a`, and \a=. Therefore we must use the latter instead of the former in the second argument to \DeclareInputText, e.g.

(from latin1.def):
     \DeclareInputText{224}{\@tabacckludge`a}

The command \@tabacckludge is defined (in ltoutenc.dtx) in such a way that \@tabacckludge' will expand to the internal form of \'. Thus it is \' that is carried around internally (the same applies to the other two accent commands).

\DeclareInputText \DeclareInputMath \IeC These commands declare the expansion of an active character. The math declaration is the usual trick with \uppercase. The text declaration is sneakier, since in text space matters. We look to see if the definition ends in a macro, by checking whether it’s \meaning ends in a space. If it does, then we add an irrelevant \IeC and braces around the definition, in order to avoid any space after the active char being gobbled up once the text is written out to an auxiliary file.

The definition should contain only robust commands (and, for correct ligatures and kerning, they must be defined via the interfaces in the fontenc package).

21 ⟨∗package⟩

22 \def\DeclareInputMath#1{%
23 \@inpenc@test
24 \bgroup
25  \uccode`\~#1%
26  \uppercase{%
27 \egroup
28  \def~%
29 }%

30 }

31 \def\DeclareInputText#1#2{%
32 \def\@tempa{#2}%
33 \def\@tempb##1 ${}%
34 \edef\@tempa{\expandafter\@tempb\meaning\@tempa$ $}%
35 \ifx\@tempa\@empty
36  \DeclareInputMath{#1}{#2}%
37 \else
38  \DeclareInputMath{#1}{\IeC{#2}}%
39 \fi

40 }

The definition of \IeC was modified not to insert a \protect unless it is needed, this means it works in \hyphenation commands, and other such delicate places. It was then further changed to never insert a \protect as one is never needed; this makes it work in even more places.

This still needs some attention.

41 \DeclareRobustCommand{\IeC}[1]{#1}
42 \def\IeC{%
43 \ifx\protect\@typeset@protect

44  \expandafter\@firstofone
45 \else
46  \noexpand\IeC
47 \fi

48 }

\inputencoding This sets the encoding to be #1. It first sets all the characters 128–255 to be active (and sets their initial definition to be \@undefined). It now also does this for some ‘low’ codes below 32, but misses out Null, control-I, control-J, control-L and control-M.

It then inputs #1.def. This command should not be used in horizontal mode since space tokens within the file will produce unwanted spaces in the output.

But it first sets up a test that produces a warning message if no suitable definitions get read.

49 \def\inputencoding#1{%
50 \gdef\@inpenc@test{\global\let\@inpenc@test\relax}%

Keyboard characters which don’t get a definition will be mapped to \@inpenc@undefined which gets the following definition:

51 \def\@inpenc@undefined{\PackageError{inputenc}%
52  {Keyboard character used is undefined\MessageBreak
53  in inputencoding #1}%
54  {You need to provide a  definition with
55  \noexpand\DeclareInputText\MessageBreak or
56  \noexpand\DeclareInputMath before using this key.}}%

Now we make all potential input characters active.

57 \ifvmode
58  \@inpenc@loop\^^A\^^H%
59  \@inpenc@loop\^^K\^^K%
60  \@inpenc@loop\^^N\^^_%
61  \@inpenc@loop\^^?\^^ff%
62  \input{#1.def}%
63 \fi
   If there have been no \DeclareInputText or \DeclareInputMath commands

read then something is amiss.

64 \ifx\@inpenc@test\relax\else
65  \PackageWarning{inputenc}%
66  {No characters defined\MessageBreak
67  by input encoding change to `#1'\MessageBreak}%
68 \fi

69 }

\@inpenc@loop Make characters #1 to #2 inclusive active and undefined.

70 \def\@inpenc@loop#1#2{%
71 \@tempcnta`#1\relax
72 \loop
73  \catcode\@tempcnta\active
74  \bgroup
75  \uccode`\~\@tempcnta
76  \uppercase{%
77  \egroup

78  \let~\@inpenc@undefined
79  }%
80 \ifnum\@tempcnta<`#2\relax
81  \advance\@tempcnta\@ne
82 \repeat}

Then for each option, we input that encoding file.

83 \DeclareOption*{\inputencoding{\CurrentOption}}
84 \ProcessOptions

85 ⟨/package⟩

4 Default definitions for characters

Some input characters map to internal functions which are not in either the T1 or OT1 font encoding. For this reason default definitions are provided in the encoding file: these will be used unless some other output encoding is used which supports those glyphs. In some cases this default definition has to be simply an error message.

Note that this works reasonably well only because the encoding files for both OT1 and T1 are loaded in the standard LaTeX format.

The name \textblacksquare is derived from the AMS symbol name since Adobe seem not to want this symbol. The default definition, as a rule, makes no claim to being a good design.

Some entries are repeated in case guards must all be on one line.

86 ⟨∗latin1 | decmulti | latin3 | latin5⟩

87 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

88 ⟨/latin1 | decmulti | latin3 | latin5⟩ 89 ⟨∗cp850 | cp852 | cp865 | cp437 | cp437de | applemac | ansinew | cp1252 | cp1250 | next⟩

90 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

91 ⟨/cp850 | cp852 | cp865 | cp437 | cp437de | applemac | ansinew | cp1252 | cp1250 | next⟩ 92 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

93 \ProvideTextCommandDefault{\textonehalf}{\ensuremath{\frac12}}

94 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 95 ⟨∗latin1 | decmulti | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

96 \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}

97 ⟨/latin1 | decmulti | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 98 ⟨∗latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

99 \ProvideTextCommandDefault{\textthreequarters}{\ensuremath{\frac34}}

100 ⟨/latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 101 ⟨∗applemac | cp850 | cp865 | cp437 | cp437de | ansinew | cp1252 | next⟩

102 \ProvideTextCommandDefault{\textflorin}{\textit{f}}

103 ⟨/applemac | cp850 | cp865 | cp437 | cp437de | ansinew | cp1252 | next⟩ 104 ⟨∗cp865 | cp437 | cp437de⟩

105 \ProvideTextCommandDefault{\textpeseta}{Pt}

106 ⟨/cp865 | cp437 | cp437de⟩ 107 ⟨∗cp850 | cp852 | cp865 | cp437 | cp437de⟩

108 \ProvideTextCommandDefault{\textblacksquare}
109 {\vrule \@width .3em \@height .4em \@depth -.1em\relax}

110 ⟨/cp850 | cp852 | cp865 | cp437 | cp437de⟩

Some commands can’t be faked, so we have them generate an error message.

111 ⟨∗latin1 | decmulti | latin5 | cp850 | cp865 | cp437 | cp437de⟩

112 \ProvideTextCommandDefault{\textcent}
113 {\TextSymbolUnavailable\textcent}
114 \ProvideTextCommandDefault{\textyen}
115 {\TextSymbolUnavailable\textyen}

116 ⟨/latin1 | decmulti | latin5 | cp850 | cp865 | cp437 | cp437de⟩ 117 ⟨∗applemac | ansinew | cp1252⟩

118 \ProvideTextCommandDefault{\textcent}
119 {\TextSymbolUnavailable\textcent}
120 \ProvideTextCommandDefault{\textyen}
121 {\TextSymbolUnavailable\textyen}

122 ⟨/applemac | ansinew | cp1252⟩ 123 ⟨∗latin1 | decmulti | latin2 | latin3 | latin5 | cp850 | cp852 | cp865⟩

124 \ProvideTextCommandDefault{\textcurrency}
125 {\TextSymbolUnavailable\textcurrency}

126 ⟨/latin1 | decmulti | latin2 | latin3 | latin5 | cp850 | cp852 | cp865⟩ 127 ⟨∗applemac | ansinew | cp1252 | cp1250⟩

128 \ProvideTextCommandDefault{\textcurrency}
129 {\TextSymbolUnavailable\textcurrency}

130 ⟨/applemac | ansinew | cp1252 | cp1250⟩ 131 ⟨∗latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | cp1250⟩

132 \ProvideTextCommandDefault{\textbrokenbar}
133 {\TextSymbolUnavailable\textbrokenbar}

134 ⟨/latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | cp1250⟩ 135 ⟨∗latin3⟩

136 \ProvideTextCommandDefault{\textmalteseH}
137 {\TextSymbolUnavailable\textmalteseH}
138 \ProvideTextCommandDefault{\textmalteseh}
139 {\TextSymbolUnavailable\textmalteseh}

140 ⟨/latin3⟩ 141 ⟨∗next⟩

142 \ProvideTextCommandDefault{\textquotesingle}
143 {\TextSymbolUnavailable\textquotesingle}

144 ⟨/next⟩ 145 ⟨∗cp1250 | ansinew | cp1252 | applemac⟩

146 \ProvideTextCommandDefault{\textperthousand}
147 {\TextSymbolUnavailable\textperthousand}

148 ⟨/cp1250 | ansinew | cp1252 | applemac⟩

Characters that are supposed to be used only in math will be defined by \providecommand because L^{A}TEX 2ε assumes that the font encoding for math fonts is static.

149 ⟨∗latin1 | decmulti | latin5 | cp850 | ansinew | cp1252 | next⟩

150 \providecommand{\mathonesuperior}{{^1}}

151 ⟨/latin1 | decmulti | latin5 | cp850 | ansinew | cp1252 | next⟩ 152 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | ansinew | cp1252 | next⟩

153 \providecommand{\maththreesuperior}{{^3}}

154 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | ansinew | cp1252 | next⟩ 155 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | cp865 | cp437 | cp437de⟩

156 \providecommand{\mathtwosuperior}{{^2}}

157 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | cp865 | cp437 | cp437de⟩ 158 ⟨∗ansinew | cp1252 | next⟩

159 \providecommand{\mathtwosuperior}{{^2}}

160 ⟨/ansinew | cp1252 | next⟩ 161 ⟨∗cp865 | cp437 | cp437de⟩

162 \providecommand{\mathnsuperior}{{^n}}

163 ⟨/cp865 | cp437 | cp437de⟩

5 The ASCII encoding

The ASCII encoding only allows characters in the range 32–127, so we only need to provide an empty .def file.

6 The ISO Latin-2 encoding

The ISO Latin-2 encoding file defines the characters in the ISO 8859-2 encoding. It was contributed by Petr Sojka (sojka@muni.cz) with small technical updates by Frank Mittelbach.

164 ⟨∗latin2⟩

165 \DeclareInputText{160}{\nobreakspace}
166 \DeclareInputText{176}{\textdegree}
167 \DeclareInputText{161}{\k A}
168 \DeclareInputText{177}{\k a}
169 \DeclareInputText{162}{\u{}}
170 \DeclareInputText{178}{\k\ }
171 \DeclareInputText{163}{\L}
172 \DeclareInputText{179}{\l}
173 \DeclareInputText{164}{\textcurrency}
174 \DeclareInputText{180}{\@tabacckludge'{}}
175 \DeclareInputText{165}{\v L}
176 \DeclareInputText{181}{\v l}
177 \DeclareInputText{166}{\@tabacckludge'S}
178 \DeclareInputText{182}{\@tabacckludge's}
179 \DeclareInputText{167}{\S}
180 \DeclareInputText{183}{\v{}}
181 \DeclareInputText{168}{\"{}}
182 \DeclareInputText{184}{\c\ }
183 \DeclareInputText{169}{\v S}
184 \DeclareInputText{185}{\v s}
185 \DeclareInputText{170}{\c S}
186 \DeclareInputText{186}{\c s}
187 \DeclareInputText{171}{\v T}
188 \DeclareInputText{187}{\v t}
189 \DeclareInputText{172}{\@tabacckludge'Z}
190 \DeclareInputText{188}{\@tabacckludge'z}
191 \DeclareInputText{173}{\-}
192 \DeclareInputText{189}{\H{}}
193 \DeclareInputText{174}{\v Z}
194 \DeclareInputText{190}{\v z}
195 \DeclareInputText{175}{\.Z}
196 \DeclareInputText{191}{\.z}
197 \DeclareInputText{192}{\@tabacckludge'R}
198 \DeclareInputText{208}{\DJ}
199 \DeclareInputText{193}{\@tabacckludge'A}
200 \DeclareInputText{209}{\@tabacckludge'N}
201 \DeclareInputText{194}{\^A}

202 \DeclareInputText{210}{\v N}
203 \DeclareInputText{195}{\u A}
204 \DeclareInputText{211}{\@tabacckludge'O}
205 \DeclareInputText{196}{\"A}
206 \DeclareInputText{212}{\^O}
207 \DeclareInputText{197}{\@tabacckludge'L}
208 \DeclareInputText{213}{\H O}
209 \DeclareInputText{198}{\@tabacckludge'C}
210 \DeclareInputText{214}{\"O}
211 \DeclareInputText{199}{\c C}
212 \DeclareInputMath{215}{\times}
213 \DeclareInputText{200}{\v C}
214 \DeclareInputText{216}{\v R}
215 \DeclareInputText{201}{\@tabacckludge'E}
216 \DeclareInputText{217}{\r U}
217 \DeclareInputText{202}{\k E}
218 \DeclareInputText{218}{\@tabacckludge'U}
219 \DeclareInputText{203}{\"E}
220 \DeclareInputText{219}{\H U}
221 \DeclareInputText{204}{\v E}
222 \DeclareInputText{220}{\"U}
223 \DeclareInputText{205}{\@tabacckludge'I}
224 \DeclareInputText{221}{\@tabacckludge'Y}
225 \DeclareInputText{206}{\^I}
226 \DeclareInputText{222}{\c T}
227 \DeclareInputText{207}{\v D}
228 \DeclareInputText{223}{\ss}
229 \DeclareInputText{224}{\@tabacckludge'r}
230 \DeclareInputText{240}{\dj}
231 \DeclareInputText{225}{\@tabacckludge'a}
232 \DeclareInputText{241}{\@tabacckludge'n}
233 \DeclareInputText{226}{\^a}
234 \DeclareInputText{242}{\v n}
235 \DeclareInputText{227}{\u a}
236 \DeclareInputText{243}{\@tabacckludge'o}
237 \DeclareInputText{228}{\"a}
238 \DeclareInputText{244}{\^o}
239 \DeclareInputText{229}{\@tabacckludge'l}
240 \DeclareInputText{245}{\H o}
241 \DeclareInputText{230}{\@tabacckludge'c}
242 \DeclareInputText{246}{\"o}
243 \DeclareInputText{231}{\c c}
244 \DeclareInputMath{247}{\div}
245 \DeclareInputText{232}{\v c}
246 \DeclareInputText{248}{\v r}
247 \DeclareInputText{233}{\@tabacckludge'e}
248 \DeclareInputText{249}{\r u}
249 \DeclareInputText{234}{\k e}
250 \DeclareInputText{250}{\@tabacckludge'u}
251 \DeclareInputText{235}{\"e}
252 \DeclareInputText{251}{\H u}
253 \DeclareInputText{236}{\v e}
254 \DeclareInputText{252}{\"u}
255 \DeclareInputText{237}{\@tabacckludge'\i}

256 \DeclareInputText{253}{\@tabacckludge'y}
257 \DeclareInputText{238}{\^\i}
258 \DeclareInputText{254}{\c t}
259 \DeclareInputText{239}{\v d}
260 \DeclareInputText{255}{\.{}}

261 ⟨/latin2⟩

7 The ISO Latin-3 encoding

The ISO Latin-3 encoding file defines the characters in the ISO 8859-3 encoding. It was contributed by by Jörg Knappen (joerg.knappen@uni-mainz.de) and adapted by Chris Rowley.

It can be used for general purpose applications in typical office environments in the following languages: Afrikaans, Catalan, English, Esperanto, French, Gallician, German, Italian, Maltese, and Turkish.

262 ⟨∗latin3⟩

263 \DeclareInputText{160}{\nobreakspace}
264 \DeclareInputText{176}{\textdegree}
265 \DeclareInputText{161}{\textmalteseH}
266 \DeclareInputText{177}{\textmalteseh}
267 \DeclareInputText{162}{\u{}}
268 \DeclareInputMath{178}{\mathtwosuperior}
269 \DeclareInputText{163}{\pounds}
270 \DeclareInputMath{179}{\maththreesuperior}
271 \DeclareInputText{164}{\textcurrency}
272 \DeclareInputText{180}{\@tabacckludge'{}}
273 % \DeclareInputText{165}{\notdef}
274 \DeclareInputMath{181}{\mu}
275 \DeclareInputText{166}{\^H}
276 % NOT: \DeclareInputText{182}{h\llap{\^{}}} % \^h would be too tall
277 \DeclareInputText{182}{\^h}
278 \DeclareInputText{167}{\S}
279 \DeclareInputText{183}{\textperiodcentered}
280 \DeclareInputText{168}{\"{}}
281 \DeclareInputText{184}{\c\ }
282 \DeclareInputText{169}{\.I}
283 \DeclareInputText{185}{\i}
284 \DeclareInputText{170}{\c S}
285 \DeclareInputText{186}{\c s}
286 \DeclareInputText{171}{\u G}
287 \DeclareInputText{187}{\u g}
288 \DeclareInputText{172}{\^J}
289 \DeclareInputText{188}{\^\j}
290 \DeclareInputText{173}{\-}
291 \DeclareInputText{189}{\textonehalf}
292 % \DeclareInputText{174}{\notdef}
293 % \DeclareInputText{190}{\notdef}
294 \DeclareInputText{175}{\.Z}
295 \DeclareInputText{191}{\.z}
296 \DeclareInputText{192}{\@tabacckludge`A}
297 % \DeclareInputText{208}{\notdef}
298 \DeclareInputText{193}{\@tabacckludge'A}

299 \DeclareInputText{209}{\~N}
300 \DeclareInputText{194}{\^A}
301 \DeclareInputText{210}{\@tabacckludge`O}
302 % \DeclareInputText{195}{\notdef}
303 \DeclareInputText{211}{\@tabacckludge'O}
304 \DeclareInputText{196}{\"A}
305 \DeclareInputText{212}{\^O}
306 \DeclareInputText{197}{\.C}
307 \DeclareInputText{213}{\.G}
308 \DeclareInputText{198}{\^C}
309 \DeclareInputText{214}{\"O}
310 \DeclareInputText{199}{\c C}
311 \DeclareInputMath{215}{\times}
312 \DeclareInputText{200}{\@tabacckludge`E}
313 \DeclareInputText{216}{\^G}
314 \DeclareInputText{201}{\@tabacckludge'E}
315 \DeclareInputText{217}{\@tabacckludge`U}
316 \DeclareInputText{202}{\^E}
317 \DeclareInputText{218}{\@tabacckludge'U}
318 \DeclareInputText{203}{\"E}
319 \DeclareInputText{219}{\^U}
320 \DeclareInputText{204}{\@tabacckludge`I}
321 \DeclareInputText{220}{\"U}
322 \DeclareInputText{205}{\@tabacckludge'I}
323 \DeclareInputText{221}{\u U}
324 \DeclareInputText{206}{\^I}
325 \DeclareInputText{222}{\^S}
326 \DeclareInputText{207}{\"I}
327 \DeclareInputText{223}{\ss}
328 \DeclareInputText{224}{\@tabacckludge`a}
329 % \DeclareInputText{240}{\notdef}
330 \DeclareInputText{225}{\@tabacckludge'a}
331 \DeclareInputText{241}{\~n}
332 \DeclareInputText{226}{\^a}
333 \DeclareInputText{242}{\@tabacckludge`o}
334 % \DeclareInputText{227}{\notdef}
335 \DeclareInputText{243}{\@tabacckludge'o}
336 \DeclareInputText{228}{\"a}
337 \DeclareInputText{244}{\^o}
338 \DeclareInputText{229}{\.c}
339 \DeclareInputText{245}{\.g}
340 \DeclareInputText{230}{\^c}
341 \DeclareInputText{246}{\"o}
342 \DeclareInputText{231}{\c c}
343 \DeclareInputMath{247}{\div}
344 \DeclareInputText{232}{\@tabacckludge`e}
345 \DeclareInputText{248}{\^g}
346 \DeclareInputText{233}{\@tabacckludge'e}
347 \DeclareInputText{249}{\@tabacckludge`u}
348 \DeclareInputText{234}{\^e}
349 \DeclareInputText{250}{\@tabacckludge'u}
350 \DeclareInputText{235}{\"e}
351 \DeclareInputText{251}{\^u}
352 \DeclareInputText{236}{\@tabacckludge`\i}

353 \DeclareInputText{252}{\"u}
354 \DeclareInputText{237}{\@tabacckludge'\i}
355 \DeclareInputText{253}{\u u}
356 \DeclareInputText{238}{\^\i}
357 \DeclareInputText{254}{\^s}
358 \DeclareInputText{239}{\"\i}
359 \DeclareInputText{255}{\.{}}

360 ⟨/latin3⟩

8 The ISO Latin-5 encoding

The ISO Latin-5 encoding file defines the characters in the ISO 8859-9 encoding, which describes Latin Alphabet No. 5. It was contributed by H. Turgut Uyar (uyar@cs.itu.edu.tr); it is used for Turkish.

361 ⟨∗latin5⟩

362 \DeclareInputText{160}{\nobreakspace}
363 \DeclareInputText{176}{\textdegree}
364 \DeclareInputText{161}{\textexclamdown}
365 \DeclareInputMath{177}{\pm}
366 \DeclareInputText{162}{\textcent}
367 \DeclareInputMath{178}{\mathtwosuperior}
368 \DeclareInputText{163}{\pounds}
369 \DeclareInputMath{179}{\maththreesuperior}
370 \DeclareInputText{164}{\textcurrency}
371 \DeclareInputText{180}{\@tabacckludge'{}}
372 \DeclareInputText{165}{\textyen}
373 \DeclareInputMath{181}{\mu}
374 \DeclareInputText{166}{\textbrokenbar}
375 \DeclareInputText{182}{\P}
376 \DeclareInputText{167}{\S}
377 \DeclareInputText{183}{\textperiodcentered}
378 \DeclareInputText{168}{\"{}}
379 \DeclareInputText{184}{\c\ }
380 \DeclareInputText{169}{\copyright}
381 \DeclareInputMath{185}{\mathonesuperior}
382 \DeclareInputText{170}{\textordfeminine}
383 \DeclareInputText{186}{\textordmasculine}
384 \DeclareInputText{171}{\guillemotleft}
385 \DeclareInputText{187}{\guillemotright}
386 \DeclareInputMath{172}{\lnot}
387 \DeclareInputText{188}{\textonequarter}
388 \DeclareInputText{173}{\-}
389 \DeclareInputText{189}{\textonehalf}
390 \DeclareInputText{174}{\textregistered}
391 \DeclareInputText{190}{\textthreequarters}
392 \DeclareInputText{175}{\@tabacckludge={}}
393 \DeclareInputText{191}{\textquestiondown}
394 \DeclareInputText{192}{\@tabacckludge`A}
395 \DeclareInputText{208}{\u G}
396 \DeclareInputText{193}{\@tabacckludge'A}
397 \DeclareInputText{209}{\~N}
398 \DeclareInputText{194}{\^A}

399 \DeclareInputText{210}{\@tabacckludge`O}
400 \DeclareInputText{195}{\~A}
401 \DeclareInputText{211}{\@tabacckludge'O}
402 \DeclareInputText{196}{\"A}
403 \DeclareInputText{212}{\^O}
404 \DeclareInputText{197}{\r A}
405 \DeclareInputText{213}{\~O}
406 \DeclareInputText{198}{\AE}
407 \DeclareInputText{214}{\"O}
408 \DeclareInputText{199}{\c C}
409 \DeclareInputMath{215}{\times}
410 \DeclareInputText{200}{\@tabacckludge`E}
411 \DeclareInputText{216}{\O}
412 \DeclareInputText{201}{\@tabacckludge'E}
413 \DeclareInputText{217}{\@tabacckludge`U}
414 \DeclareInputText{202}{\^E}
415 \DeclareInputText{218}{\@tabacckludge'U}
416 \DeclareInputText{203}{\"E}
417 \DeclareInputText{219}{\^U}
418 \DeclareInputText{204}{\@tabacckludge`I}
419 \DeclareInputText{220}{\"U}
420 \DeclareInputText{205}{\@tabacckludge'I}
421 \DeclareInputText{221}{\.I}
422 \DeclareInputText{206}{\^I}
423 \DeclareInputText{222}{\c S}
424 \DeclareInputText{207}{\"I}
425 \DeclareInputText{223}{\ss}
426 \DeclareInputText{224}{\@tabacckludge`a}
427 \DeclareInputText{240}{\u g}
428 \DeclareInputText{225}{\@tabacckludge'a}
429 \DeclareInputText{241}{\~n}
430 \DeclareInputText{226}{\^a}
431 \DeclareInputText{242}{\@tabacckludge`o}
432 \DeclareInputText{227}{\~a}
433 \DeclareInputText{243}{\@tabacckludge'o}
434 \DeclareInputText{228}{\"a}
435 \DeclareInputText{244}{\^o}
436 \DeclareInputText{229}{\r a}
437 \DeclareInputText{245}{\~o}
438 \DeclareInputText{230}{\ae}
439 \DeclareInputText{246}{\"o}
440 \DeclareInputText{231}{\c c}
441 \DeclareInputMath{247}{\div}
442 \DeclareInputText{232}{\@tabacckludge`e}
443 \DeclareInputText{248}{\o}
444 \DeclareInputText{233}{\@tabacckludge'e}
445 \DeclareInputText{249}{\@tabacckludge`u}
446 \DeclareInputText{234}{\^e}
447 \DeclareInputText{250}{\@tabacckludge'u}
448 \DeclareInputText{235}{\"e}
449 \DeclareInputText{251}{\^u}
450 \DeclareInputText{236}{\@tabacckludge`\i}
451 \DeclareInputText{252}{\"u}
452 \DeclareInputText{237}{\@tabacckludge'\i}

453 \DeclareInputText{253}{\i }
454 \DeclareInputText{238}{\^\i}
455 \DeclareInputText{254}{\c s}
456 \DeclareInputText{239}{\"\i}
457 \DeclareInputText{255}{\"y}

458 ⟨/latin5⟩

9 DEC Multinational Character Set

The DECMultinational character set, used by the OpenVMS operating system, is slightly different from the ISO Latin 1 character set.

Reference: Digital Equipment Corporation VT330/VT340 Programmer Reference Manual, Volume 1: Text Programming, page 22.

This encoding was provided by M.Y. Chartoire IPNL-IN2P3

 m.chartoire@ipnl.in2p3.fr

459 ⟨∗decmulti⟩

460 \DeclareInputText{176}{\textdegree}
461 \DeclareInputText{161}{\textexclamdown}
462 \DeclareInputMath{177}{\pm}
463 \DeclareInputText{162}{\textcent}
464 \DeclareInputMath{178}{\mathtwosuperior}
465 \DeclareInputText{163}{\pounds}
466 \DeclareInputMath{179}{\maththreesuperior}
467 \DeclareInputText{165}{\textyen}
468 \DeclareInputMath{181}{\mu}
469 \DeclareInputText{182}{\P}
470 \DeclareInputText{167}{\S}
471 \DeclareInputText{183}{\textperiodcentered}
472 \DeclareInputText{168}{\textcurrency}
473 \DeclareInputText{169}{\copyright}
474 \DeclareInputMath{185}{\mathonesuperior}
475 \DeclareInputText{170}{\textordfeminine}
476 \DeclareInputText{186}{\textordmasculine}
477 \DeclareInputText{171}{\guillemotleft}
478 \DeclareInputText{187}{\guillemotright}
479 \DeclareInputText{188}{\textonequarter}
480 \DeclareInputText{189}{\textonehalf}
481 \DeclareInputText{191}{\textquestiondown}
482 \DeclareInputText{192}{\@tabacckludge`A}
483 \DeclareInputText{193}{\@tabacckludge'A}
484 \DeclareInputText{209}{\~N}
485 \DeclareInputText{194}{\^A}
486 \DeclareInputText{210}{\@tabacckludge`O}
487 \DeclareInputText{195}{\~A}
488 \DeclareInputText{211}{\@tabacckludge'O}
489 \DeclareInputText{196}{\"A}
490 \DeclareInputText{212}{\^O}
491 \DeclareInputText{197}{\r A}
492 \DeclareInputText{213}{\~O}
493 \DeclareInputText{198}{\AE}
494 \DeclareInputText{214}{\"O}
495 \DeclareInputText{199}{\c C}

496 \DeclareInputText{215}{\OE}
497 \DeclareInputText{200}{\@tabacckludge`E}
498 \DeclareInputText{216}{\O}
499 \DeclareInputText{201}{\@tabacckludge'E}
500 \DeclareInputText{217}{\@tabacckludge`U}
501 \DeclareInputText{202}{\^E}
502 \DeclareInputText{218}{\@tabacckludge'U}
503 \DeclareInputText{203}{\"E}
504 \DeclareInputText{219}{\^U}
505 \DeclareInputText{204}{\@tabacckludge`I}
506 \DeclareInputText{220}{\"U}
507 \DeclareInputText{205}{\@tabacckludge'I}
508 \DeclareInputText{221}{\"Y}
509 \DeclareInputText{206}{\^I}
510 \DeclareInputText{207}{\"I}
511 \DeclareInputText{223}{\ss}
512 \DeclareInputText{224}{\@tabacckludge`a}
513 \DeclareInputText{225}{\@tabacckludge'a}
514 \DeclareInputText{241}{\~n}
515 \DeclareInputText{226}{\^a}
516 \DeclareInputText{242}{\@tabacckludge`o}
517 \DeclareInputText{227}{\~a}
518 \DeclareInputText{243}{\@tabacckludge'o}
519 \DeclareInputText{228}{\"a}
520 \DeclareInputText{244}{\^o}
521 \DeclareInputText{229}{\r a}
522 \DeclareInputText{245}{\~o}
523 \DeclareInputText{230}{\ae}
524 \DeclareInputText{246}{\"o}
525 \DeclareInputText{231}{\c c}
526 \DeclareInputText{247}{\oe}
527 \DeclareInputText{232}{\@tabacckludge`e}
528 \DeclareInputText{248}{\o}
529 \DeclareInputText{233}{\@tabacckludge'e}
530 \DeclareInputText{249}{\@tabacckludge`u}
531 \DeclareInputText{234}{\^e}
532 \DeclareInputText{250}{\@tabacckludge'u}
533 \DeclareInputText{235}{\"e}
534 \DeclareInputText{251}{\^u}
535 \DeclareInputText{236}{\@tabacckludge`\i}
536 \DeclareInputText{252}{\"u}
537 \DeclareInputText{237}{\@tabacckludge'\i}
538 \DeclareInputText{253}{\"y}
539 \DeclareInputText{238}{\^\i}
540 \DeclareInputText{239}{\"\i}

541 ⟨/decmulti⟩

10 The IBM code page 850

This input encoding was contributed by Timo Knuutila (knuutila@cs.utu.fi), and edited by Christian Bartels (ii140ba@vm1.rz.rwth-aachen.de).

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

542 ⟨∗cp850⟩

543 \DeclareInputText{128}{\c C}
544 \DeclareInputText{144}{\@tabacckludge'E}
545 \DeclareInputText{129}{\"u}
546 \DeclareInputText{145}{\ae}
547 \DeclareInputText{130}{\@tabacckludge'e}
548 \DeclareInputText{146}{\AE}
549 \DeclareInputText{131}{\^a}
550 \DeclareInputText{147}{\^o}
551 \DeclareInputText{132}{\"a}
552 \DeclareInputText{148}{\"o}
553 \DeclareInputText{133}{\@tabacckludge`a}
554 \DeclareInputText{149}{\@tabacckludge`o}
555 \DeclareInputText{134}{\r a}
556 \DeclareInputText{150}{\^u}
557 \DeclareInputText{135}{\c c}
558 \DeclareInputText{151}{\@tabacckludge`u}
559 \DeclareInputText{136}{\^e}
560 \DeclareInputText{152}{\"y}
561 \DeclareInputText{137}{\"e}
562 \DeclareInputText{153}{\"O}
563 \DeclareInputText{138}{\@tabacckludge`e}
564 \DeclareInputText{154}{\"U}
565 \DeclareInputText{139}{\"\i}
566 \DeclareInputText{155}{\o}
567 \DeclareInputText{140}{\^\i}
568 \DeclareInputText{156}{\pounds}
569 \DeclareInputText{141}{\@tabacckludge`\i}
570 \DeclareInputText{157}{\O}
571 \DeclareInputText{142}{\"A}
572 \DeclareInputMath{158}{\times}
573 \DeclareInputText{143}{\r A}
574 \DeclareInputText{159}{\textflorin}
575 \DeclareInputText{160}{\@tabacckludge'a}
576 \DeclareInputText{161}{\@tabacckludge'\i}
577 \DeclareInputText{162}{\@tabacckludge'o}
578 \DeclareInputText{163}{\@tabacckludge'u}
579 \DeclareInputText{164}{\~n}               %% "B0-"B4: DG
580 \DeclareInputText{165}{\~N}
581 \DeclareInputText{181}{\@tabacckludge'A}
582 \DeclareInputText{166}{\textordfeminine}
583 \DeclareInputText{182}{\^A}
584 \DeclareInputText{167}{\textordmasculine}
585 \DeclareInputText{183}{\@tabacckludge`A}
586 \DeclareInputText{168}{\textquestiondown}
587 \DeclareInputText{184}{\copyright}
588 \DeclareInputText{169}{\textregistered}
589 \DeclareInputMath{170}{\lnot}
590 \DeclareInputText{171}{\textonehalf}
591 \DeclareInputText{172}{\textonequarter}   %% "B9-"BC: DG
592 \DeclareInputText{173}{\textexclamdown}
593 \DeclareInputText{189}{\textcent}

594 \DeclareInputText{174}{\guillemotleft}
595 \DeclareInputText{190}{\textyen}
596 \DeclareInputText{175}{\guillemotright}
597 % "BF-"C5: DG
598 \DeclareInputText{208}{\dh}
599 \DeclareInputText{209}{\DH}
600 \DeclareInputText{210}{\^E}
601 \DeclareInputText{211}{\"E}
602 \DeclareInputText{212}{\@tabacckludge`E}
603 \DeclareInputText{213}{\i}
604 \DeclareInputText{198}{\~a}
605 \DeclareInputText{214}{\@tabacckludge'I}
606 \DeclareInputText{199}{\~A}
607 \DeclareInputText{215}{\^I}
608 \DeclareInputText{216}{\"I}
609 % "C8-"CE: DG
610 % "D9-"DC: DG
611 \DeclareInputText{221}{\textbrokenbar}
612 \DeclareInputText{222}{\@tabacckludge`I}
613 \DeclareInputText{207}{\textcurrency}  %% "DF: DG
614 \DeclareInputText{224}{\@tabacckludge'O}
615 \DeclareInputText{240}{\textendash}
616 \DeclareInputText{225}{\ss}
617 \DeclareInputMath{241}{\pm}
618 \DeclareInputText{226}{\^O}
619 % "F2: DG (not double underline, or equals?)
620 \DeclareInputText{227}{\@tabacckludge`O}
621 \DeclareInputText{243}{\textthreequarters}
622 \DeclareInputText{228}{\~o}
623 \DeclareInputText{244}{\P}
624 \DeclareInputText{229}{\~O}
625 \DeclareInputText{245}{\S}
626 \DeclareInputMath{230}{\mu}
627 \DeclareInputMath{246}{\div}
628 \DeclareInputText{231}{\th}
629 \DeclareInputText{247}{\c\ }
630 \DeclareInputText{232}{\TH}
631 \DeclareInputText{248}{\textdegree}
632 \DeclareInputText{233}{\@tabacckludge'U}
633 \DeclareInputText{249}{\"{}}
634 \DeclareInputText{234}{\^U}
635 \DeclareInputText{250}{\textperiodcentered}
636 \DeclareInputText{235}{\@tabacckludge`U}
637 \DeclareInputMath{251}{\mathonesuperior}
638 \DeclareInputText{236}{\@tabacckludge'y}
639 \DeclareInputMath{252}{\maththreesuperior}
640 \DeclareInputText{237}{\@tabacckludge'Y}
641 \DeclareInputMath{253}{\mathtwosuperior}
642 \DeclareInputText{238}{\@tabacckludge={}}
643 \DeclareInputText{254}{\textblacksquare} % right name?
644 \DeclareInputText{239}{\@tabacckludge'{}}
645 \DeclareInputText{255}{\nobreakspace}

646 ⟨/cp850⟩

11 The IBM code page 852

This input encoding was contributed by Petr Sojka (sojka@Muni.cz).

647 ⟨∗cp852⟩

648 \DeclareInputText{128}{\c C}
649 \DeclareInputText{144}{\@tabacckludge'E}
650 \DeclareInputText{129}{\"u}
651 \DeclareInputText{145}{\@tabacckludge'L}
652 \DeclareInputText{130}{\@tabacckludge'e}
653 \DeclareInputText{146}{\@tabacckludge'l}
654 \DeclareInputText{131}{\^a}
655 \DeclareInputText{147}{\^o}
656 \DeclareInputText{132}{\"a}
657 \DeclareInputText{148}{\"o}
658 \DeclareInputText{133}{\r u}
659 \DeclareInputText{149}{\v L}
660 \DeclareInputText{134}{\@tabacckludge'c}
661 \DeclareInputText{150}{\v l}
662 \DeclareInputText{135}{\c c}
663 \DeclareInputText{151}{\@tabacckludge'S}
664 \DeclareInputText{136}{\l}
665 \DeclareInputText{152}{\@tabacckludge's}
666 \DeclareInputText{137}{\"e}
667 \DeclareInputText{153}{\"O}
668 \DeclareInputText{138}{\H O}
669 \DeclareInputText{154}{\"U}
670 \DeclareInputText{139}{\H o}
671 \DeclareInputText{155}{\v T}
672 \DeclareInputText{140}{\^\i}
673 \DeclareInputText{156}{\v t}
674 \DeclareInputText{141}{\@tabacckludge'Z}
675 \DeclareInputText{157}{\L}
676 \DeclareInputText{142}{\"A}
677 \DeclareInputMath{158}{\times}
678 \DeclareInputText{143}{\@tabacckludge'C}
679 \DeclareInputText{159}{\v c}
680 \DeclareInputText{160}{\@tabacckludge'a}
681 \DeclareInputText{161}{\@tabacckludge'\i}
682 \DeclareInputText{162}{\@tabacckludge'o}
683 \DeclareInputText{163}{\@tabacckludge'u}
684 \DeclareInputText{164}{\k A}
685 \DeclareInputText{165}{\k a}
686 \DeclareInputText{166}{\v Z}
687 \DeclareInputText{167}{\v z}
688 \DeclareInputText{168}{\k E}
689 \DeclareInputText{169}{\k e}
690 \DeclareInputText{170}{\nobreakspace}  % or empty??
691 \DeclareInputText{171}{\@tabacckludge'z}
692 \DeclareInputText{172}{\v C}
693 \DeclareInputText{173}{\c s}
694 \DeclareInputText{174}{\guillemotleft}
695 \DeclareInputText{175}{\guillemotright}
696 \DeclareInputText{181}{\@tabacckludge'A}

697 \DeclareInputText{182}{\^A}
698 \DeclareInputText{183}{\v E}
699 \DeclareInputText{184}{\c S}
700 \DeclareInputText{189}{\.Z}
701 \DeclareInputText{190}{\.z}
702 \DeclareInputText{198}{\u A}
703 \DeclareInputText{199}{\u a}
704 \DeclareInputText{207}{\textcurrency}
705 \DeclareInputText{208}{\dj}
706 \DeclareInputText{209}{\DJ}
707 \DeclareInputText{210}{\v D}
708 \DeclareInputText{211}{\"E}
709 \DeclareInputText{212}{\v d} % d caron
710 \DeclareInputText{213}{\v N}
711 \DeclareInputText{214}{\@tabacckludge'I}
712 \DeclareInputText{215}{\^I}
713 \DeclareInputText{216}{\v e}
714 \DeclareInputText{221}{\c T}
715 \DeclareInputText{222}{\r U}
716 \DeclareInputText{224}{\@tabacckludge'O}
717 \DeclareInputText{240}{\textendash}
718 \DeclareInputText{225}{\ss}
719 \DeclareInputText{241}{\H{}}
720 \DeclareInputText{226}{\^O}
721 \DeclareInputText{242}{\k\ }
722 \DeclareInputText{227}{\@tabacckludge'N}
723 \DeclareInputText{243}{\v{}}
724 \DeclareInputText{228}{\@tabacckludge'n}
725 \DeclareInputText{244}{\u{}}
726 \DeclareInputText{229}{\v n}
727 \DeclareInputText{245}{\S}
728 \DeclareInputText{230}{\v S}
729 \DeclareInputMath{246}{\div}
730 \DeclareInputText{231}{\v s}
731 \DeclareInputText{247}{\c\ }
732 \DeclareInputText{232}{\@tabacckludge'R}
733 \DeclareInputText{248}{\textdegree}
734 \DeclareInputText{233}{\@tabacckludge'U}
735 \DeclareInputText{249}{\"{}}
736 \DeclareInputText{234}{\@tabacckludge'r}
737 \DeclareInputText{250}{\.{}}
738 \DeclareInputText{235}{\H U}
739 \DeclareInputText{251}{\H u}
740 \DeclareInputText{236}{\@tabacckludge'y}
741 \DeclareInputText{252}{\v R}
742 \DeclareInputText{237}{\@tabacckludge'Y}
743 \DeclareInputText{253}{\v r}
744 \DeclareInputText{238}{\c t}
745 \DeclareInputText{254}{\textblacksquare} % right name?
746 \DeclareInputText{239}{\@tabacckludge'{}}
747 \DeclareInputText{255}{\nobreakspace}

748 ⟨/cp852⟩

12 The IBM code pages 437 and 865

This input encoding is based on work by

 Volker Kunert (volker@numsun1.mathematik.uni-halle.de)
 and bontus@al6000.physik.uni-siegen.de.

The changes for cp865 are based on work by Søren Sandmann (sandmann@ daimi.aau.dk), with thanks to them all.

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

Unfortunately, in cp437 there is no agreement as to whether slot E1 should be ‘β’ or ‘ß’, so we provide two variants, one (cp437) with ‘β’ and one (cp437de) with ‘ß’.

749 ⟨∗cp437 | cp437de | cp865⟩

750 \DeclareInputText{128}{{\c C}}
751 \DeclareInputText{129}{\"u}
752 \DeclareInputText{130}{\@tabacckludge'e}
753 \DeclareInputText{131}{\^a}
754 \DeclareInputText{132}{\"a}
755 \DeclareInputText{133}{\@tabacckludge`a}
756 \DeclareInputText{134}{\r a}
757 \DeclareInputText{135}{{\c c}}
758 \DeclareInputText{136}{\^e}
759 \DeclareInputText{137}{\"e}
760 \DeclareInputText{138}{\@tabacckludge`e}
761 \DeclareInputText{139}{\"\i}
762 \DeclareInputText{140}{\^\i}
763 \DeclareInputText{141}{\@tabacckludge`\i}
764 \DeclareInputText{142}{\"A}
765 \DeclareInputText{143}{\r A}
766 \DeclareInputText{144}{\@tabacckludge'E}
767 \DeclareInputText{145}{\ae}
768 \DeclareInputText{146}{\AE}
769 \DeclareInputText{147}{\^o}
770 \DeclareInputText{148}{\"o}
771 \DeclareInputText{149}{\@tabacckludge`o}
772 \DeclareInputText{150}{\^u}
773 \DeclareInputText{151}{\@tabacckludge`u}
774 \DeclareInputText{152}{\"y}
775 \DeclareInputText{153}{\"O}
776 \DeclareInputText{154}{\"U}

777 ⟨/cp437 | cp437de | cp865⟩ 778 ⟨∗cp437 | cp437de⟩

779 \DeclareInputText{155}{\textcent}
780 \DeclareInputText{156}{\pounds}
781 \DeclareInputText{157}{\textyen}

782 ⟨/cp437 | cp437de⟩ 783 ⟨∗cp865⟩

784 \DeclareInputText{155}{\o}
785 \DeclareInputText{156}{\pounds}
786 \DeclareInputText{157}{\O}

787 ⟨/cp865⟩ 788 ⟨∗cp437 | cp437de | cp865⟩

789 \DeclareInputText{158}{\textpeseta} % Pt

790 \DeclareInputText{159}{\textflorin}
791 \DeclareInputText{160}{\@tabacckludge'a} % 160
792 \DeclareInputText{161}{\@tabacckludge'\i}
793 \DeclareInputText{162}{\@tabacckludge'o}
794 \DeclareInputText{163}{\@tabacckludge'u}
795 \DeclareInputText{164}{\~n}
796 \DeclareInputText{165}{\~N}
797 \DeclareInputText{166}{\textordfeminine}
798 \DeclareInputText{167}{\textordmasculine}
799 \DeclareInputText{168}{\textquestiondown}
800 %\DeclareInputText{169}{} % left upper corner
801 \DeclareInputMath{170}{\lnot}
802 \DeclareInputText{171}{\textonehalf}
803 \DeclareInputText{172}{\textonequarter}
804 \DeclareInputText{173}{\textexclamdown}
805 \DeclareInputText{174}{\guillemotleft}
806 ⟨cp437 | cp437de⟩\DeclareInputText{175}{\guillemotright}
807 ⟨cp865⟩\DeclareInputText{175}{\textcurrency}
808 %\DeclareInputText{176}{\textlightgraybox}
809 %\DeclareInputText{177}{\textgraybox}
810 %\DeclareInputText{178}{\textdarkgraybox}
811 %\DeclareInputMath{179}{} % vertical bar
812 %\DeclareInputText{180}{} % vertical bar with branch to left
813 %\DeclareInputText{181}{} % vertical bar with double branch to left
814 %\DeclareInputText{182}{} % double bar with single branch to left
815 %\DeclareInputText{183}{} % graphic
816 %\DeclareInputText{184}{} % graphic
817 %\DeclareInputMath{185}{} % vertical double bar with branch to left
818 %\DeclareInputMath{186}{} % vertical double bar
819 %\DeclareInputText{187}{} % double upper right corner
820 %\DeclareInputMath{188}{} % double lower right corner
821 %\DeclareInputText{189}{} % graphic
822 %\DeclareInputText{190}{} % graphic
823 %\DeclareInputMath{191}{\ensuremath{\rceil}}
824 %\DeclareInputMath{192}{\ensuremath{\lfloor}}
825 %\DeclareInputText{193}{} % dash with branch up
826 %\DeclareInputText{194}{} % dash with branch down
827 %\DeclareInputText{195}{} % vertical bar with branch to right
828 %\DeclareInputText{196}{} % horizontal bar, not endash
829 %\DeclareInputText{197}{} % vertical bar crossed with dash
830 %\DeclareInputText{198}{} % graphic
831 %\DeclareInputText{199}{} % graphic
832 %\DeclareInputText{200}{} % double lower left corner
833 %\DeclareInputText{201}{} % double upper left corner
834 %\DeclareInputText{202}{} % double dash with branch up
835 %\DeclareInputText{203}{} % double dash with branch down
836 %\DeclareInputText{204}{} % double bar with branch right
837 %\DeclareInputText{205}{=} % double dash
838 %\DeclareInputText{206}{} % double bar crossing double dash
839 %\DeclareInputText{207}{} % graphic
840 %\DeclareInputMath{208}{}
841 %\DeclareInputText{209}{}
842 %\DeclareInputText{210}{}

843 %\DeclareInputText{211}{}
844 %\DeclareInputText{212}{}
845 %\DeclareInputText{213}{}
846 %\DeclareInputText{214}{}
847 %\DeclareInputText{215}{}
848 %\DeclareInputText{216}{}
849 %\DeclareInputMath{217}{\ensuremath{\rfloor}} % lower right corner
850 %\DeclareInputMath{218}{\ensuremath{\lceil}}  % upper left corner
851 %\DeclareInputText{219}{} % black box
852 %\DeclareInputText{220}{} % lower half of black box
853 %\DeclareInputText{221}{} % left bar
854 %\DeclareInputText{222}{} % right bar
855 %\DeclareInputText{223}{} % upper half of black box
856 \DeclareInputMath{224}{\alpha}

857 ⟨/cp437 | cp437de | cp865⟩

858 ⟨cp437 | cp865⟩\DeclareInputMath{225}{\beta}
859 ⟨cp437de⟩\DeclareInputText{225}{\ss}

860 ⟨∗cp437 | cp437de | cp865⟩

861 \DeclareInputMath{226}{\Gamma}
862 \DeclareInputMath{227}{\pi}
863 \DeclareInputMath{228}{\Sigma}
864 \DeclareInputMath{229}{\sigma}
865 \DeclareInputMath{230}{\mu}
866 \DeclareInputMath{231}{\gamma}
867 \DeclareInputMath{232}{\Phi}
868 \DeclareInputMath{233}{\theta}
869 \DeclareInputMath{234}{\Omega}
870 \DeclareInputMath{235}{\delta}
871 \DeclareInputMath{236}{\infty}
872 \DeclareInputMath{237}{\phi}
873 \DeclareInputMath{238}{\varepsilon}
874 \DeclareInputMath{239}{\cap}
875 \DeclareInputMath{240}{\equiv}
876 \DeclareInputMath{241}{\pm}
877 \DeclareInputMath{242}{\ge}
878 \DeclareInputMath{243}{\le}
879 %\DeclareInputMath{244}{}   % upper part of integral sign
880 %\DeclareInputMath{245}{}   % lower part of integral sign
881 \DeclareInputMath{246}{\div}
882 \DeclareInputMath{247}{\approx}
883 \DeclareInputText{248}{\textdegree}
884 \DeclareInputText{249}{\textperiodcentered}
885 \DeclareInputText{250}{\textbullet}
886 \DeclareInputMath{251}{\surd}
887 \DeclareInputMath{252}{\mathnsuperior}
888 \DeclareInputMath{253}{\mathtwosuperior}
889 \DeclareInputText{254}{\textblacksquare} % right name?
890 \DeclareInputText{255}{\nobreakspace}

891 ⟨/cp437 | cp437de | cp865⟩

13 The Macintosh encoding

This input encoding was contributed by Constantin Kahn (kahn@math.toronto.edu), with minor modifications by Alan Jeffrey.

892 ⟨∗applemac⟩

893 \DeclareInputText{128}{\"A}
894 \DeclareInputText{129}{\r A}
895 \DeclareInputText{130}{\c C}
896 \DeclareInputText{131}{\@tabacckludge'E}
897 \DeclareInputText{132}{\~N}
898 \DeclareInputText{133}{\"O}
899 \DeclareInputText{134}{\"U}
900 \DeclareInputText{135}{\@tabacckludge'a}
901 \DeclareInputText{136}{\@tabacckludge`a}
902 \DeclareInputText{137}{\^a}
903 \DeclareInputText{138}{\"a}
904 \DeclareInputText{139}{\~a}
905 \DeclareInputText{140}{\r a}
906 \DeclareInputText{141}{\c c}
907 \DeclareInputText{142}{\@tabacckludge'e}
908 \DeclareInputText{143}{\@tabacckludge`e}
909 \DeclareInputText{144}{\^e}
910 \DeclareInputText{145}{\"e}
911 \DeclareInputText{146}{\@tabacckludge'\i}
912 \DeclareInputText{147}{\@tabacckludge`\i}
913 \DeclareInputText{148}{\^\i}
914 \DeclareInputText{149}{\"\i}
915 \DeclareInputText{150}{\~n}
916 \DeclareInputText{151}{\@tabacckludge'o}
917 \DeclareInputText{152}{\@tabacckludge`o}
918 \DeclareInputText{153}{\^o}
919 \DeclareInputText{154}{\"o}
920 \DeclareInputText{155}{\~o}
921 \DeclareInputText{156}{\@tabacckludge'u}
922 \DeclareInputText{157}{\@tabacckludge`u}
923 \DeclareInputText{158}{\^u}
924 \DeclareInputText{159}{\"u}
925 \DeclareInputText{160}{\dag}
926 \DeclareInputText{161}{\textdegree}
927 \DeclareInputText{162}{\textcent}
928 \DeclareInputText{163}{\pounds}
929 \DeclareInputText{164}{\S}
930 \DeclareInputText{165}{\textbullet}
931 \DeclareInputText{166}{\P}
932 \DeclareInputText{167}{\ss}
933 \DeclareInputText{168}{\textregistered}
934 \DeclareInputText{169}{\copyright}
935 \DeclareInputText{170}{\texttrademark}
936 \DeclareInputText{171}{\@tabacckludge'{}}
937 \DeclareInputText{172}{\"{}}
938 \DeclareInputMath{173}{\neq}
939 \DeclareInputText{174}{\AE}
940 \DeclareInputText{175}{\O}

941 \DeclareInputMath{176}{\infty}
942 \DeclareInputMath{177}{\pm}
943 \DeclareInputMath{178}{\leq}
944 \DeclareInputMath{179}{\geq}
945 \DeclareInputText{180}{\textyen}
946 \DeclareInputMath{181}{\mu}
947 \DeclareInputMath{182}{\partial}
948 \DeclareInputMath{183}{\Sigma}
949 \DeclareInputMath{184}{\Pi}
950 \DeclareInputMath{185}{\pi}
951 \DeclareInputMath{186}{\int}
952 \DeclareInputText{187}{\textordfeminine}
953 \DeclareInputText{188}{\textordmasculine}
954 \DeclareInputMath{189}{\Omega}
955 \DeclareInputText{190}{\ae}
956 \DeclareInputText{191}{\o}
957 \DeclareInputText{192}{\textquestiondown}
958 \DeclareInputText{193}{\textexclamdown}
959 \DeclareInputMath{194}{\lnot}
960 \DeclareInputMath{195}{\surd}
961 \DeclareInputText{196}{\textflorin}
962 \DeclareInputMath{197}{\approx}
963 \DeclareInputMath{198}{\Delta}
964 \DeclareInputText{199}{\guillemotleft}
965 \DeclareInputText{200}{\guillemotright}
966 \DeclareInputText{201}{\dots}
967 \DeclareInputText{202}{\nobreakspace}
968 \DeclareInputText{203}{\@tabacckludge`A}
969 \DeclareInputText{204}{\~A}
970 \DeclareInputText{205}{\~O}
971 \DeclareInputText{206}{\OE}
972 \DeclareInputText{207}{\oe}
973 \DeclareInputText{208}{\textendash}
974 \DeclareInputText{209}{\textemdash}
975 \DeclareInputText{210}{\textquotedblleft}
976 \DeclareInputText{211}{\textquotedblright}
977 \DeclareInputText{212}{\textquoteleft}
978 \DeclareInputText{213}{\textquoteright}
979 \DeclareInputMath{214}{\div}
980 \DeclareInputMath{215}{\diamond}
981 \DeclareInputText{216}{\"y}
982 \DeclareInputText{217}{\"Y}
983 \DeclareInputMath{218}{/}
984 \DeclareInputText{219}{\textcurrency}
985 \DeclareInputText{220}{\guilsinglleft}
986 \DeclareInputText{221}{\guilsinglright}
987 \DeclareInputText{222}{fi}
988 \DeclareInputText{223}{fl}
989 \DeclareInputText{224}{\ddag}
990 \DeclareInputText{225}{\textperiodcentered}
991 \DeclareInputText{226}{\quotesinglbase}
992 \DeclareInputText{227}{\quotedblbase}
993 \DeclareInputText{228}{\textperthousand}

 994 \DeclareInputText{229}{\^A}
 995 \DeclareInputText{230}{\^E}
 996 \DeclareInputText{231}{\@tabacckludge'A}
 997 \DeclareInputText{232}{\"E}
 998 \DeclareInputText{233}{\@tabacckludge`E}
 999 \DeclareInputText{234}{\@tabacckludge'I}
1000 \DeclareInputText{235}{\^I}
1001 \DeclareInputText{236}{\"I}
1002 \DeclareInputText{237}{\@tabacckludge`I}
1003 \DeclareInputText{238}{\@tabacckludge'O}
1004 \DeclareInputText{239}{\^O}
1005 \DeclareInputText{240}{\textapplelogo}
1006 \DeclareInputText{241}{\@tabacckludge`O}
1007 \DeclareInputText{242}{\@tabacckludge'U}
1008 \DeclareInputText{243}{\^U}
1009 \DeclareInputText{244}{\@tabacckludge`U}
1010 \DeclareInputText{245}{\i}
1011 \DeclareInputText{246}{\^{}}
1012 \DeclareInputText{247}{\~{}}
1013 \DeclareInputText{248}{\@tabacckludge={}}
1014 \DeclareInputText{249}{\u{}}
1015 \DeclareInputText{250}{\.{}}
1016 \DeclareInputText{251}{\r{}}
1017 \DeclareInputText{252}{\c\ }
1018 \DeclareInputText{253}{\H{}}
1019 \DeclareInputText{254}{\k\ }
1020 \DeclareInputText{255}{\v{}}

1021 ⟨/applemac⟩

14 The Next encoding

This input encoding is based on work by Stefan Ried (stef@theo-phys.uni-essen.de

  and Holger Uhr (huhr@uni-paderborn.de).

1022 ⟨∗next⟩

1023 \DeclareInputText{128}{\nobreakspace}
1024 \DeclareInputText{129}{\@tabacckludge`A}
1025 \DeclareInputText{130}{\@tabacckludge'A}
1026 \DeclareInputText{131}{\^A}
1027 \DeclareInputText{132}{\~A}
1028 \DeclareInputText{133}{\"A}
1029 \DeclareInputText{134}{\r A}
1030 \DeclareInputText{135}{\c C}
1031 \DeclareInputText{136}{\@tabacckludge`E}
1032 \DeclareInputText{137}{\@tabacckludge'E}
1033 \DeclareInputText{138}{\^E}
1034 \DeclareInputText{139}{\"E}
1035 \DeclareInputText{140}{\@tabacckludge`I}
1036 \DeclareInputText{141}{\@tabacckludge'I}
1037 \DeclareInputText{142}{\^I}
1038 \DeclareInputText{143}{\"I}
1039 %\DeclareInputText{144}{?}
1040 \DeclareInputText{145}{\~N}
1041 \DeclareInputText{146}{\@tabacckludge`O}

1042 \DeclareInputText{147}{\@tabacckludge'O}
1043 \DeclareInputText{148}{\^O}
1044 \DeclareInputText{149}{\~O}
1045 \DeclareInputText{150}{\"O}
1046 \DeclareInputText{151}{\@tabacckludge`U}
1047 \DeclareInputText{152}{\@tabacckludge'U}
1048 \DeclareInputText{153}{\^U}
1049 \DeclareInputText{154}{\"U}
1050 \DeclareInputText{155}{\@tabacckludge'Y}
1051 %\DeclareInputText{156}{?}
1052 \DeclareInputMath{157}{\mu}
1053 \DeclareInputMath{158}{\times}
1054 \DeclareInputMath{159}{\div}
1055 \DeclareInputText{160}{\copyright}
1056 \DeclareInputText{161}{\textexclamdown}
1057 %\DeclareInputText{162}{?}
1058 \DeclareInputText{163}{\pounds}
1059 \DeclareInputMath{164}{/}
1060 %\DeclareInputText{165}{?}
1061 \DeclareInputText{166}{\textflorin}
1062 \DeclareInputText{167}{\S}
1063 %\DeclareInputText{168}{?}
1064 \DeclareInputText{169}{\textquotesingle}
1065 \DeclareInputText{170}{\textquotedblleft}
1066 %\DeclareInputText{171}{?}
1067 %\DeclareInputText{172}{?}
1068 %\DeclareInputText{173}{?}
1069 \DeclareInputText{174}{fi}
1070 \DeclareInputText{175}{fl}
1071 %\DeclareInputText{176}{?}
1072 \DeclareInputText{177}{\textendash}
1073 \DeclareInputText{178}{\dag}
1074 \DeclareInputText{179}{\ddag}
1075 \DeclareInputText{180}{\textperiodcentered}
1076 \DeclareInputText{181}{\textbar}
1077 \DeclareInputText{182}{\P}
1078 \DeclareInputText{183}{\textbullet}
1079 \DeclareInputText{184}{\quotesinglbase}
1080 \DeclareInputText{185}{\quotedblbase}
1081 \DeclareInputText{186}{\textquotedblright}
1082 %\DeclareInputText{187}{?}
1083 \DeclareInputText{188}{\dots}
1084 %\DeclareInputText{189}{?}
1085 \DeclareInputMath{190}{\lnot}
1086 \DeclareInputText{191}{\textquestiondown}
1087 \DeclareInputMath{192}{\mathonesuperior}
1088 \DeclareInputText{193}{\@tabacckludge`{}}
1089 \DeclareInputText{194}{\@tabacckludge'{}}
1090 \DeclareInputText{195}{\^{}}
1091 \DeclareInputText{196}{\~{}}
1092 \DeclareInputText{197}{\@tabacckludge={}}
1093 \DeclareInputText{198}{\u{}}
1094 \DeclareInputText{199}{\.{}}

1095 \DeclareInputText{200}{\"{}}
1096 \DeclareInputMath{201}{\mathtwosuperior}
1097 \DeclareInputText{202}{\textdegree}
1098 \DeclareInputText{203}{\c\ }
1099 \DeclareInputMath{204}{\maththreesuperior}
1100 \DeclareInputText{205}{\H{}}
1101 %\DeclareInputText{206}{?}
1102 \DeclareInputText{207}{\v{}}
1103 \DeclareInputText{208}{\textemdash}
1104 \DeclareInputMath{209}{\pm}
1105 \DeclareInputText{210}{\textonequarter}
1106 \DeclareInputText{211}{\textonehalf}
1107 \DeclareInputText{212}{\textthreequarters}
1108 \DeclareInputText{213}{\@tabacckludge`a}
1109 \DeclareInputText{214}{\@tabacckludge'a}
1110 \DeclareInputText{215}{\^a}
1111 \DeclareInputText{216}{\~a}
1112 \DeclareInputText{217}{\"a}
1113 \DeclareInputText{218}{\r a}
1114 \DeclareInputText{219}{\c c}
1115 \DeclareInputText{220}{\@tabacckludge`e}
1116 \DeclareInputText{221}{\@tabacckludge'e}
1117 \DeclareInputText{222}{\^e}
1118 \DeclareInputText{223}{\"e}
1119 \DeclareInputText{224}{\@tabacckludge`\i}
1120 \DeclareInputText{225}{\AE}
1121 \DeclareInputText{226}{\@tabacckludge'\i}
1122 \DeclareInputText{227}{\textordfeminine}
1123 \DeclareInputText{228}{\^\i}
1124 \DeclareInputText{229}{\"\i}
1125 %\DeclareInputText{230}{?}
1126 \DeclareInputText{231}{\~n}
1127 \DeclareInputText{232}{\L}
1128 \DeclareInputText{233}{\O}
1129 \DeclareInputText{234}{\OE}
1130 \DeclareInputText{235}{\textordmasculine}
1131 \DeclareInputText{236}{\@tabacckludge`o}
1132 \DeclareInputText{237}{\@tabacckludge'o}
1133 \DeclareInputText{238}{\^o}
1134 \DeclareInputText{239}{\~o}
1135 \DeclareInputText{240}{\"o}
1136 \DeclareInputText{241}{\ae}
1137 \DeclareInputText{242}{\@tabacckludge`u}
1138 \DeclareInputText{243}{\@tabacckludge'u}
1139 \DeclareInputText{244}{\^u}
1140 \DeclareInputText{245}{\i}
1141 \DeclareInputText{246}{\"u}
1142 \DeclareInputText{247}{\@tabacckludge'y}
1143 \DeclareInputText{248}{\l}
1144 \DeclareInputText{249}{\o}
1145 \DeclareInputText{250}{\oe}
1146 \DeclareInputText{251}{\ss}
1147 %\DeclareInputText{252}{?}
1148 \DeclareInputText{253}{\"y}

1149 ⟨/next⟩

15 The MS Windows ANSI encoding cp1252, and the ISO Latin-1 encoding

The MS Windows ANSI, cp 1252, input encoding was contributed by Berthold

  K.P. Horn (bkph@ai.mit.edu).

It has two very different names so the same code produces two files. These encoding files both define the characters in the MS Windows 3.1 ANSI encoding (Western Europe), also known as code page 1252, which is based on ISO Latin-1 but has important additions in the 128–159 range.

Designed for: Danish, Dutch, English, Finnish, French, German, Icelandic, Italian, Norwegian, Portuguese, Spanish, and Swedish.

Note: Windows ANSI — like Macintosh standard roman encoding — has quotesingle at 39, and grave at 96 — which is here ignored.

The ISO Latin-1 encoding file defines only the restricted range of characters available in the ISO 8859-1 encoding.

1150 ⟨∗ansinew | cp1252⟩

1151 \DeclareInputText{130}{\quotesinglbase}
1152 \DeclareInputText{131}{\textflorin}
1153 \DeclareInputText{132}{\quotedblbase}
1154 \DeclareInputText{133}{\dots}
1155 \DeclareInputText{134}{\dag}
1156 \DeclareInputText{135}{\ddag}
1157 \DeclareInputText{136}{\^{}}
1158 \DeclareInputText{137}{\textperthousand}
1159 \DeclareInputText{138}{\v S}
1160 \DeclareInputText{139}{\guilsinglleft}
1161 \DeclareInputText{140}{\OE}
1162 \DeclareInputText{145}{\textquoteleft}
1163 \DeclareInputText{146}{\textquoteright}
1164 \DeclareInputText{147}{\textquotedblleft}
1165 \DeclareInputText{148}{\textquotedblright}
1166 \DeclareInputText{149}{\textbullet}
1167 \DeclareInputText{150}{\textendash}
1168 \DeclareInputText{151}{\textemdash}
1169 \DeclareInputText{152}{\~{}}
1170 \DeclareInputText{153}{\texttrademark}
1171 \DeclareInputText{154}{\v s}
1172 \DeclareInputText{155}{\guilsinglright}
1173 \DeclareInputText{156}{\oe}
1174 \DeclareInputText{159}{\"Y}

1175 ⟨/ansinew | cp1252⟩

1176 ⟨∗ansinew | cp1252 | latin1⟩

1177 \DeclareInputText{160}{\nobreakspace}
1178 \DeclareInputText{176}{\textdegree}
1179 \DeclareInputText{161}{\textexclamdown}
1180 \DeclareInputMath{177}{\pm}
1181 \DeclareInputText{162}{\textcent}
1182 \DeclareInputMath{178}{\mathtwosuperior}
1183 \DeclareInputText{163}{\pounds}

1184 \DeclareInputMath{179}{\maththreesuperior}
1185 \DeclareInputText{164}{\textcurrency}
1186 \DeclareInputText{180}{\@tabacckludge'{}}
1187 \DeclareInputText{165}{\textyen}
1188 \DeclareInputMath{181}{\mu}
1189 \DeclareInputText{166}{\textbrokenbar}
1190 \DeclareInputText{182}{\P}
1191 \DeclareInputText{167}{\S}
1192 \DeclareInputText{183}{\textperiodcentered}
1193 \DeclareInputText{168}{\"{}}
1194 \DeclareInputText{184}{\c\ }
1195 \DeclareInputText{169}{\copyright}
1196 \DeclareInputMath{185}{\mathonesuperior}
1197 \DeclareInputText{170}{\textordfeminine}
1198 \DeclareInputText{186}{\textordmasculine}
1199 \DeclareInputText{171}{\guillemotleft}
1200 \DeclareInputText{187}{\guillemotright}
1201 \DeclareInputMath{172}{\lnot}
1202 \DeclareInputText{188}{\textonequarter}
1203 \DeclareInputText{173}{\-}
1204 \DeclareInputText{189}{\textonehalf}
1205 \DeclareInputText{174}{\textregistered}
1206 \DeclareInputText{190}{\textthreequarters}
1207 \DeclareInputText{175}{\@tabacckludge={}}
1208 \DeclareInputText{191}{\textquestiondown}
1209 \DeclareInputText{192}{\@tabacckludge`A}
1210 \DeclareInputText{208}{\DH}
1211 \DeclareInputText{193}{\@tabacckludge'A}
1212 \DeclareInputText{209}{\~N}
1213 \DeclareInputText{194}{\^A}
1214 \DeclareInputText{210}{\@tabacckludge`O}
1215 \DeclareInputText{195}{\~A}
1216 \DeclareInputText{211}{\@tabacckludge'O}
1217 \DeclareInputText{196}{\"A}
1218 \DeclareInputText{212}{\^O}
1219 \DeclareInputText{197}{\r A}
1220 \DeclareInputText{213}{\~O}
1221 \DeclareInputText{198}{\AE}
1222 \DeclareInputText{214}{\"O}
1223 \DeclareInputText{199}{\c C}
1224 \DeclareInputMath{215}{\times}
1225 \DeclareInputText{200}{\@tabacckludge`E}
1226 \DeclareInputText{216}{\O}
1227 \DeclareInputText{201}{\@tabacckludge'E}
1228 \DeclareInputText{217}{\@tabacckludge`U}
1229 \DeclareInputText{202}{\^E}
1230 \DeclareInputText{218}{\@tabacckludge'U}
1231 \DeclareInputText{203}{\"E}
1232 \DeclareInputText{219}{\^U}
1233 \DeclareInputText{204}{\@tabacckludge`I}
1234 \DeclareInputText{220}{\"U}
1235 \DeclareInputText{205}{\@tabacckludge'I}
1236 \DeclareInputText{221}{\@tabacckludge'Y}
1237 \DeclareInputText{206}{\^I}

1238 \DeclareInputText{222}{\TH}
1239 \DeclareInputText{207}{\"I}
1240 \DeclareInputText{223}{\ss}
1241 \DeclareInputText{224}{\@tabacckludge`a}
1242 \DeclareInputText{240}{\dh}
1243 \DeclareInputText{225}{\@tabacckludge'a}
1244 \DeclareInputText{241}{\~n}
1245 \DeclareInputText{226}{\^a}
1246 \DeclareInputText{242}{\@tabacckludge`o}
1247 \DeclareInputText{227}{\~a}
1248 \DeclareInputText{243}{\@tabacckludge'o}
1249 \DeclareInputText{228}{\"a}
1250 \DeclareInputText{244}{\^o}
1251 \DeclareInputText{229}{\r a}
1252 \DeclareInputText{245}{\~o}
1253 \DeclareInputText{230}{\ae}
1254 \DeclareInputText{246}{\"o}
1255 \DeclareInputText{231}{\c c}
1256 \DeclareInputMath{247}{\div}
1257 \DeclareInputText{232}{\@tabacckludge`e}
1258 \DeclareInputText{248}{\o}
1259 \DeclareInputText{233}{\@tabacckludge'e}
1260 \DeclareInputText{249}{\@tabacckludge`u}
1261 \DeclareInputText{234}{\^e}
1262 \DeclareInputText{250}{\@tabacckludge'u}
1263 \DeclareInputText{235}{\"e}
1264 \DeclareInputText{251}{\^u}
1265 \DeclareInputText{236}{\@tabacckludge`\i}
1266 \DeclareInputText{252}{\"u}
1267 \DeclareInputText{237}{\@tabacckludge'\i}
1268 \DeclareInputText{253}{\@tabacckludge'y}
1269 \DeclareInputText{238}{\^\i}
1270 \DeclareInputText{254}{\th}
1271 \DeclareInputText{239}{\"\i}
1272 \DeclareInputText{255}{\"y}

1273 ⟨/ansinew | cp1252 | latin1⟩

16 The MS Windows encoding cp1250

This is an MS Windows encoding for Central and Eastern Europe, known as Code Page 1250; it was provided by Marcin Woliński wolinski@melkor.mimuw.edu.pl and edited by Chris Rowley (see v0.94 below) to make it consistent with other files.

1274 ⟨∗cp1250⟩

1275 %\DeclareInputText{128}{} % not used
1276 %\DeclareInputText{129}{} % n/u
1277 \DeclareInputText{130}{\quotesinglbase}
1278 %\DeclareInputText{131}{} % n/u
1279 \DeclareInputText{132}{\quotedblbase}
1280 \DeclareInputText{133}{\dots}
1281 \DeclareInputText{134}{\dag}
1282 \DeclareInputText{135}{\ddag}

1283 %\DeclareInputText{136}{} % n/u
1284 \DeclareInputText{137}{\textperthousand}
1285 \DeclareInputText{138}{\v S}
1286 \DeclareInputText{139}{\guilsinglleft}
1287 \DeclareInputText{140}{\@tabacckludge'S}
1288 \DeclareInputText{141}{\v T}
1289 \DeclareInputText{142}{\v Z}
1290 \DeclareInputText{143}{\@tabacckludge'Z}
1291 %\DeclareInputText{144}{} % n/u
1292 \DeclareInputText{145}{\textquoteleft}
1293 \DeclareInputText{146}{\textquoteright}
1294 \DeclareInputText{147}{\textquotedblleft}
1295 \DeclareInputText{148}{\textquotedblright}
1296 \DeclareInputText{149}{\textbullet}
1297 \DeclareInputText{150}{\textendash}
1298 \DeclareInputText{151}{\textemdash}
1299 %\DeclareInputText{152}{} % n/u
1300 \DeclareInputText{153}{\texttrademark}
1301 \DeclareInputText{154}{\v s}
1302 \DeclareInputText{155}{\guilsinglright}
1303 \DeclareInputText{156}{\@tabacckludge's}
1304 \DeclareInputText{157}{\v t} % t caron (t') ?
1305 \DeclareInputText{158}{\v z}
1306 \DeclareInputText{159}{\@tabacckludge'z}
1307 \DeclareInputText{160}{\nobreakspace}
1308 \DeclareInputText{161}{\v{}}
1309 \DeclareInputText{162}{\u{}}
1310 \DeclareInputText{163}{\L}
1311 \DeclareInputText{164}{\textcurrency}
1312 \DeclareInputText{165}{\k A}
1313 \DeclareInputText{166}{\textbrokenbar}
1314 \DeclareInputText{167}{\S}
1315 \DeclareInputText{168}{\"{}}
1316 \DeclareInputText{169}{\copyright}
1317 \DeclareInputText{170}{\c S}
1318 \DeclareInputText{171}{\guillemotleft}
1319 \DeclareInputMath{172}{\lnot}
1320 \DeclareInputText{173}{\-} % soft hyphen
1321 \DeclareInputText{174}{\textregistered}
1322 \DeclareInputText{175}{\.Z}
1323 \DeclareInputText{176}{\textdegree} % degree sign
1324 \DeclareInputMath{177}{\pm} % plus-minus
1325 \DeclareInputText{178}{\k\ }
1326 \DeclareInputText{179}{\l}
1327 \DeclareInputText{180}{\@tabacckludge'{}}
1328 \DeclareInputMath{181}{\mu} % micro sign
1329 \DeclareInputText{182}{\P}
1330 \DeclareInputText{183}{\textperiodcentered}
1331 \DeclareInputText{184}{\c\ }
1332 \DeclareInputText{185}{\k a}
1333 \DeclareInputText{186}{\c s}
1334 \DeclareInputText{187}{\guillemotright}
1335 \DeclareInputText{188}{\v L} % L caron (L') ?
1336 \DeclareInputText{189}{\H{}}

1337 \DeclareInputText{190}{\v l} % l caron (l') ?
1338 \DeclareInputText{191}{\. z}
1339 \DeclareInputText{192}{\@tabacckludge'R}
1340 \DeclareInputText{193}{\@tabacckludge'A}
1341 \DeclareInputText{194}{\^A}
1342 \DeclareInputText{195}{\u A}
1343 \DeclareInputText{196}{\"A}
1344 \DeclareInputText{197}{\@tabacckludge'L}
1345 \DeclareInputText{198}{\@tabacckludge'C}
1346 \DeclareInputText{199}{\c C}
1347 \DeclareInputText{200}{\v C}
1348 \DeclareInputText{201}{\@tabacckludge'E}
1349 \DeclareInputText{202}{\k E}
1350 \DeclareInputText{203}{\"E}
1351 \DeclareInputText{204}{\v E}
1352 \DeclareInputText{205}{\@tabacckludge'I}
1353 \DeclareInputText{206}{\^I}
1354 \DeclareInputText{207}{\v D}
1355 \DeclareInputText{208}{\DJ} % D stroke
1356 \DeclareInputText{209}{\@tabacckludge'N}
1357 \DeclareInputText{210}{\v N}
1358 \DeclareInputText{211}{\@tabacckludge'O}
1359 \DeclareInputText{212}{\^O}
1360 \DeclareInputText{213}{\H O}
1361 \DeclareInputText{214}{\"O}
1362 \DeclareInputMath{215}{\times}
1363 \DeclareInputText{216}{\v R}
1364 \DeclareInputText{217}{\r U}
1365 \DeclareInputText{218}{\@tabacckludge'U}
1366 \DeclareInputText{219}{\H U}
1367 \DeclareInputText{220}{\" U}
1368 \DeclareInputText{221}{\@tabacckludge'Y}
1369 \DeclareInputText{222}{\c T}
1370 \DeclareInputText{223}{\ss}
1371 \DeclareInputText{224}{\@tabacckludge'r}
1372 \DeclareInputText{225}{\@tabacckludge'a}
1373 \DeclareInputText{226}{\^a}
1374 \DeclareInputText{227}{\u a}
1375 \DeclareInputText{228}{\" a}
1376 \DeclareInputText{229}{\@tabacckludge'l}
1377 \DeclareInputText{230}{\@tabacckludge'c}
1378 \DeclareInputText{231}{\c c}
1379 \DeclareInputText{232}{\v c}
1380 \DeclareInputText{233}{\@tabacckludge'e}
1381 \DeclareInputText{234}{\k e}
1382 \DeclareInputText{235}{\"e}
1383 \DeclareInputText{236}{\v e}
1384 \DeclareInputText{237}{\@tabacckludge'\i}
1385 \DeclareInputText{238}{\^\i}
1386 \DeclareInputText{239}{\v d} % d caron (d') ?
1387 \DeclareInputText{240}{\dj} % d stroke
1388 \DeclareInputText{241}{\@tabacckludge'n}
1389 \DeclareInputText{242}{\v n}

1390 \DeclareInputText{243}{\@tabacckludge'o}
1391 \DeclareInputText{244}{\^o}
1392 \DeclareInputText{245}{\H o}
1393 \DeclareInputText{246}{\"o}
1394 \DeclareInputMath{247}{\div}
1395 \DeclareInputText{248}{\v r}
1396 \DeclareInputText{249}{\r u}
1397 \DeclareInputText{250}{\@tabacckludge'u}
1398 \DeclareInputText{251}{\H u}
1399 \DeclareInputText{252}{\"u}
1400 \DeclareInputText{253}{\@tabacckludge'y}
1401 \DeclareInputText{254}{\c t}
1402 \DeclareInputText{255}{\.{}}

1403 ⟨/cp1250⟩

Finally, we reset the category code of the at sign at the end of all .def files.

1404 ⟨−package⟩\makeatother